OpenStudy (anonymous):

P=(V^2R)/(R+r)^2 V, R are constants; r is a variable Find dP/dr

OpenStudy (anonymous):

P' = V^2R * d(1/(R+r)^2)/dr P' = V^2R * d((R+r)^(-2))/dr P' = V^2R * ( -2*(R+r)^(-3))

OpenStudy (amistre64):

\[P=(V^2R).(R+r)^{-2}\] P' = (V^2 R) -2(R+r)^-3 is what I get

OpenStudy (amistre64):

[V^2 R]' = 0 [(R+r)^-2]' = -2(R+r)^-3

OpenStudy (anonymous):

According to the software, -2(R+r)^-3 is not the correct answer.

OpenStudy (amistre64):

thats becasue it has a negative exponent. I just assumed you had some background in these things

OpenStudy (amistre64):

\[P' = -\frac{(V^2 R)}{2(R+r)^3}\]

OpenStudy (anonymous):

The software automatically converts the negative exponents, so there was no need for me to do anything with it. Your most recent answer is still coming up wrong.

OpenStudy (amistre64):

thats becasue you stated that V was a constant as well

OpenStudy (amistre64):

V get derived with respecr to 'r' then as well since it aint a constant

OpenStudy (anonymous):

V is a constant. That's what the professor told us in class.

OpenStudy (amistre64):

your screenshot does not say V is a constant; have you been using V as a constant in this? or are you just assuming that some problem that your prof used where V was a constant applies to all Vs that you come across?

OpenStudy (anonymous):

He told us in class that, for this specific problem, V is a constant. He said the authors of the book failed to include that piece of information in the problem.

OpenStudy (amistre64):

Then if you can find where I have an error at, then let me know. Cuase I cant see it :)

OpenStudy (anonymous):

Me neither. This problem is driving me nuts. :P

OpenStudy (amistre64):

i think I do; put the 2 up top; I dragged it along and it should stay up top

OpenStudy (amistre64):

\[-2(R+r)^{-3}=\frac{-2}{(R+r)^3}\]

OpenStudy (anonymous):

Yep. That was it. :D

OpenStudy (anonymous):

But it did make me keep the \[V ^{2}R\]

OpenStudy (amistre64):

lol ..... accursed programs :)

OpenStudy (anonymous):

Indeed. Thank you for all your help though!